OpenStudy (anonymous):

if 0 is less than or equal to t which is less than 2pi solve cos t + 1=2sin^2t any idea?

OpenStudy (anonymous):

Do you have options?

OpenStudy (anonymous):

No all on our own

OpenStudy (anonymous):

cos(t) + 1 = 2 sin^2 t substitute , sin^2 t = 1 - cos^2 t from the identity sin^2 t + cos^2 t = 1 cos(t) + 1 = 2 ( 1 - cos^2 t ) distribute cos (t) + 1 = 2 - 2 cos^2 t 2 cos^2 (t) + cos (t) + 1 - 2 = 0 2 cos^2 (t) + cos (t) -1 = 0 factor (2 cos t - 1 ) ( cos t + 1 ) = 0

OpenStudy (anonymous):

@SurpriseBreeze i think you forgot to divide by two tho cuz look at number for equation

OpenStudy (anonymous):

four**

OpenStudy (anonymous):

I just googled it... sorry .-. I like your elephant! :D

OpenStudy (anonymous):

@SurpriseBreeze So is your answer right tho? and haha thanks